A string with free boundary caused by stationary waves. The distance of four adjacent nodes is 90cm. If the speed of wave propagation is 45m/s. the frequency of the wave is...


1
Expert's answer
2021-06-08T09:35:48-0400

"l=90cm=0.9m"

"n=4"

"\lambda=\frac{l}{n}=\frac{0.9}{4}=0.225"

"v=45\frac{m}{s}"

"\lambda=\frac{v}{f}"

"f=\frac{v}{\lambda}=\frac{45}{0.225}=200\ Hz"

"\text{Answer: }f=200\ Hz"